OpenStudy (anonymous):

2x 11 3/7 = 4 2/5 x 3 1/11=

OpenStudy (hihi67):

(a^2+ab)*(2+ab+ab^2)

OpenStudy (phi):

are these 2 different problems? \[2 \times 11\frac{3}{7}= \] \[ 4 \frac{2}{5} \times 3 \frac{1}{11}= \]

OpenStudy (phi):

One way to do this, is change the fraction to an improper fraction then multiply.

OpenStudy (phi):

another way is to use the distributive law. 11 3/7 is the same as \[ (11+\frac{3}{7}) \] so \[ 2(11+\frac{3}{7}) = 2 \times 11+2 \times \frac{3}{7}\]